OpenStudy (anonymous):

verify cos^2x-sin^2x=2cos^2x-1

Parth (parthkohli):

There are two ways to do this. One: substitute \(\sin^2(x) \) with \(1 - \cos^2(x)\). Second, notice that the left-hand side is equal to \(\cos(2x)\). I prefer the first one.